[figure: 9.7 cm]

.

MAGNETIC AND ELECTRIC STUDIES OF FeSb_{2−x}Cr_{x}S_{4}

A.I. Djabbarov Institute of Physics Academy of Sciences of Azerbaijan, Huseyn Javid St.33, 370143 Baku, Azerbaijan

Magnetic semiconductors of FeSb_{2−x}Cr_{x}S_{4} 0≤ x ≤2 were prepared and the lattice parameters evaluated as a function of x at room temperature, show deviation from Vegard’s law. This was attributed to the presence of magnetovolume effect and charge transfer. Magnetization measurements were performed in the temperature range 77-300K. The results suggest that the system undergoes a gradual transition from antiferromagnetic to ferrimagnetic with increasing x. This has been explained based on the 3d band model. Furthermore, a monotonic decrease of the Curie temperature with x was observed. Electrical resistivity studies were carried out in the temperature range 4.2-300K. An anomaly was found in the vicinity of the Curie (Néel) temperature for x=2 which has been attributed to the existence of the long-range nature of the critical fluctuations due to short-range spin ordering.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : A.I.Djabbarov

Address for correspondence : Institute of Physics Academy of Sciences of Azerbaijan, Huseyn Javid St.33, 370143 Baku, Azerbaijan

Email address : physic@lan.ab.az